OpenStudy (haleyelizabeth2017):

Solve the system \(|y| \ge 2\) and \(|x| \le 1\) by graphing. Please no direct answers :)

jimthompson5910 (jim_thompson5910):

hint: \[\Large |y| \ge 2 \text{ is the same as } -2 \le y \le 2\]

OpenStudy (haleyelizabeth2017):

Oh!

OpenStudy (haleyelizabeth2017):

Thank you for clarifying that!

jimthompson5910 (jim_thompson5910):

oh sorry, I mixed up the signs I was focusing on x

jimthompson5910 (jim_thompson5910):

\[\Large |x| \le 1 \text{ is the same as } -1 \le x \le 1\]

jimthompson5910 (jim_thompson5910):

Rule: \[\Large |x| \le k \text{ is equivalent to } -k \le x \le k\] where k is any positive number

OpenStudy (haleyelizabeth2017):

Ah hah

jimthompson5910 (jim_thompson5910):

Rule: \[\Large |x| \ge k \text{ is equivalent to } x \ge k \text{ or } x \le -k\] where k is any positive number

OpenStudy (zzr0ck3r):

they want you to graph it

jimthompson5910 (jim_thompson5910):

using that second rule, \[\Large |y| \ge 2\] breaks down into \[\Large y \ge 2 \text{ or } y \le -2\]

jimthompson5910 (jim_thompson5910):

So you need to graph the following \[\Large -2 \le x \le 2, \ y \ge 2, \ y \le -2\]

OpenStudy (haleyelizabeth2017):

Okay so the solution is where they intersect/overlap?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (haleyelizabeth2017):

Thank you so very much!

jimthompson5910 (jim_thompson5910):

looks good so far